3 qt. of a sugar solution was mixed with 4 qt. of a 4% sugar solution to make a 40% sugar solution. Find the percent concentration of the first solution.
Answer by ptaylor(2198)

OK
Let x=the percent sugar in the 3qt solution
Now we know that the amount of pure sugar in the 3qt solution (3x) plus the amount of pure sugar in the 4qt solution (0.04*4) has to equal the amount of pure sugar in the final solution (7*0.40), so our equation to solve is:
3x+0.04*4=0.40*7 simplify
3x+0.16=2.8 subtract 0.16 from each side
3x+0.16-0.16=2.8-0.16 collect like terms
3x=2.64 divide each side by 3
x=0.88 or 88%----the percent sugar in the 3qt solution

CK
3*0.88+4*0.04=0.40*7
2.64+0.16=2.80
2.80=2.80
Hope this helps----ptaylor